# EUROPEAN ORGANIZATION FOR NUCLEAR RESEARCH (CERN) CERN-EP-2024-125 LHCb-PAPER-2024-007

### v1

@@
LHCb collaboration†††Authors are listed at the end of this letter.
The branching fraction ratios of and decays are measured with respect to their muonic counterparts, using a data sample corresponding to an integrated luminosity of 2.0 fb-1 collected by the LHCb experiment in proton-proton collisions at . The reconstructed final states are formed by combining mesons with candidates, where the is reconstructed via the decay. The results are
@@
Submitted to Phys. Rev. Lett.
© 2024 CERN for the benefit of the LHCb collaboration. CC BY 4.0 licence.
1 Introduction
@@
In the Standard Model (SM), the transition proceeds via a tree-level boson exchange.111The inclusion of charge conjugate processes in this Letter is implied throughout. The SM decay amplitude is therefore large, resulting in large branching fractions and copious signal yields. Despite this, transitions are highly sensitive to new physics as the process involves three third-generation fermions [1, 2, 3, 4]. This system therefore represents one of the best potential sources of contributions from new physics mediators that couple preferentially to the third generation [5, 6]. Furthermore, being a charged-current semileptonic transition, non-factorisable long-distance contributions are strongly suppressed. Uncertainties due to the strong interaction can be reduced by forming lepton universality ratios, where a semitauonic transition is normalised to the corresponding semimuonic transition, . In the SM, these ratios only deviate from unity due to the different lepton masses.
@@
The experimental challenge in analysing this system is the presence of multiple neutrinos in the final state, mandating a partial reconstruction of both the -lepton and -hadron decays. This partial reconstruction leaves the signal highly susceptible to a large number of backgrounds, which must be carefully controlled. Previous analyses of the system include measurements of the -lepton and charm-hadron polarisation [7, 8] and the lepton universality ratios and , which are formed from the branching fraction ratios . A combination [9] of the measurements [10, 11, 12, 13, 14, 15, 16, 17] differs by approximately three standard deviations from the SM predictions [18, 19, 20, 21, 22, 23, 24, 25, 26], indicating a potential enhancement of the semitauonic transition rate. Additional measurements of this process are therefore highly motivated.
@@
This Letter describes the first simultaneous measurement of the lepton universality ratios at the LHCb experiment, using the decays , and . Using isospin symmetry, this can be interpreted as a measurement of . The analysis strategy is similar to the existing measurement performed with the decay [17], where a three-dimensional template fit to the squared four-momentum transferred to the lepton pair (), the energy of the muon in the meson rest frame () and the squared invariant mass missing from the visible system () is performed to separate the signal and the normalisation from the backgrounds. The contribution from the first excited states ( and ) to the ground state is significantly reduced in this measurement compared to the case in Ref. [17]. The meson in the decay is not reconstructed, due to the large background from neutral objects. This measurement uses data corresponding to 2.0 collected during the period 2015–2016 in 13 TeV collisions.
2 Detector and dataset
@@
Simulated events are used to optimise the signal selection, model the signal and backgrounds, and calculate the relative efficiency between the signal and normalisation channels. In the simulation, collisions are generated using Pythia [30] with a specific LHCb configuration [32]. Decays of hadronic particles are described by EvtGen [33], in which final-state radiation is generated using Photos [34]. The interaction of the generated particles with the detector, and its response, are implemented using the Geant4 toolkit [35] as described in Ref. [37]. For this Letter, a new fast simulation technique is used whereby the photon propagation in the Ring Imaging CHerenkov (RICH) detectors, shower development in the calorimeters and activity in the muon stations are not simulated. This tracker-only simulation is a factor of eight faster and has a 40% smaller event size than full simulation. A smaller sample of simulated events which includes the full set of interactions (full simulation) is used for validation and to develop specific tools such as those to reject partially reconstructed backgrounds with an additional neutral pion or photon in the final state.
3 Selection
@@
At the hardware stage, events are required to be triggered by activity unrelated to the signal candidate or by high transverse energy signals in the HCAL due to the decay products. In the software stage, one- or two-track combinations of the decay products are required to pass a multivariate selection based on the transverse momentum and impact parameter with respect to an associated primary vertex (PV). A subsequent software level selects a displaced three-track combination with a invariant mass compatible with the mass and have large transverse momentum. The candidate is combined with a muon with a loose vertex-quality requirement. The muon is required to have a momentum above 3 to ensure that it reaches the muon stations. Particle identification (PID) requirements are applied on the final state particles, based on information from the RICH and muon systems.
@@
A centralised preselection, which follows the final trigger stage selection, is additionally applied. The candidate is required to have a significant impact parameter with respect to any PV, which reduces background from charm decays originating directly from the collision. A neural network classifier [38] is also applied to all final-state particles to reject track candidates which have more than 30% of hits that do not originate from the same source (fake tracks).
@@
Background for which the candidate is an accidental track combination from different sources is rejected by employing a Boosted Decision Tree (BDT) [39, 40] implemented in the TMVA toolkit [41]. The BDT is trained to distinguish simulated signal decays from background, which is modelled using data selected with invariant mass outside the range 1840–1900. The BDT features include kinematic and geometric information about the candidate as well as track quality variables for the decay products. The BDT requirement is optimised by maximising the metric , where is the expected number of signal decays in simulation and is the amount of background extrapolated from the mass sidebands in the data. This requirement is 97% efficient for the signal and rejects 35% of the background.
@@
Although the decay , where , is a signal channel in this analysis, the selection is optimised for the decay and therefore an isolation requirement is imposed based on neutral calorimeter objects, typically produced in partially-reconstructed background decays. A dedicated BDT is trained to discriminate between and decays, which is referred to as neutral isolation. This BDT uses the presence of neutral calorimeter objects in a cone around the direction and reconstructed candidates that have an invariant mass consistent with the known mass [43].
@@
Finally, a BDT-based isolation tool for charged tracks [44], referred to as track isolation, is used to classify all additional tracks in the event based on whether they are consistent with originating from the -decay vertex. The three tracks that are most likely to originate from the signal candidate are used to control partially reconstructed backgrounds.
@@
A maximum-likelihood fit to the mass distribution of the selected candidates is performed, where the signal is parameterised by an Hypatia function [45] and the background by an exponential function. The fit results in 3.1M candidates is used to define weights in the sPlot technique [46], such that background where the combination does not originate from a meson is subtracted in the determination of the signal yield.
@@ 4 Corrections to simulation
Several corrections are applied to improve the accuracy of the simulation. Firstly, the trigger and neutral isolation selection must be emulated for the tracker-only simulation. The hardware-level trigger is emulated by reproducing the two ways in which the events can be triggered. The efficiency for triggering via other particles in the event, which accounts for around 90% of the candidates, is measured using decays in data as a function of the kinematics. The emulation of the efficiency for one of the decay products to directly trigger through a high cluster in the HCAL is determined with fully simulated signal decays as a function of the kinematics of the decay products.
The offline selection provides an excellent description of the software trigger, with the exception of the PID response, which is corrected for separately. The efficiency of the neutral isolation BDT is emulated using full simulation by analysing the kinematic correlation of the candidate and generated photons in its vicinity, and employing a BDT based algorithm [47] to account for correlations with the fit variables.
@@
The efficiencies of the PID requirements on the reconstructed final-state particles are determined using control samples [48] as a function of track kinematics and event multiplicity. They are also used to determine misidentification probabilities, which are crucial to estimate the background where the muon candidate is a misidentified track. The tracking efficiency is determined using decays in data [49], as a function of the track kinematics. The kinematics of the meson as well as the event multiplicity are corrected using decays.
@@
Finally, a weight-based calibration procedure is designed to correct a comprehensive set of geometric and kinematic variables associated with the final-state particles as well as the Dalitz variables of the decay. In addition to the properties of the candidate, the kinematics of the three additional tracks found by the track-isolation algorithm are also corrected. This correction procedure compares the selected candidates in data with the simulation in the region , where the sample is enriched in the normalisation channels. The result of this procedure depends on the relative abundance of the different decays assumed in the simulation, which is estimated by an initial fit to the data sample. The corrected simulation is consistent with the data in the distributions of all considered variables.
@@
The corrected simulation is used to determine the relative efficiency between the signal and normalisation channels. This efficiency ratio is found to be 0.57 for the mode and 0.58 for the mode. These ratios are smaller than unity since the relatively low momentum muon from the lepton has a small detection efficiency, and the lifetime results in a poorly defined vertex. Systematic uncertainties on this ratio are discussed in Sect. 6 and are subdominant compared to those affecting the signal yield determination.
5 Signal yield determination
@@
The tauonic signal decay is only 2% of the selected sample, meaning that control over the background is essential to obtain a precise and accurate measurement. A three-dimensional binned-likelihood fit to , and is therefore employed to fully exploit the kinematic differences in the signal and background channels. Each contribution to the fit is modelled using a template histogram derived either from simulation or from specifically selected control samples in data.
@@
Uncertainties in the modelling of the components are incorporated into the fit as bin-by-bin correlated variations in the templates. These variations are implemented via a polynomial interpolation [50]. Exceptions to this are variations in the form factors, which are performed using the HAMMER tool [51] as implemented in Ref. [52]. The HAMMER tool allows for a fast, but exact, variation of a template as a function of the decay-model parameters. This variation is important to allow the form factors of both the signal and normalisation channels to vary as even the constraints derived from precise lattice calculations [53, 26, 54, 55] can have larger uncertainties than those obtained from the fit. The form factors for the and modes are parameterised according to the BGL formalism [56] with multivariate Gaussian constraints applied according to the predictions reported in Refs. [57, 26].
@@
The background originating from decays is less well-understood compared to those involving the lower mass states, despite a recent measurement by Belle [58]. This is because the measurements are less precise and the narrow-width assumption is less well-motivated due to the significant natural width of the states. The form-factor models employed for the decays are those described in Ref. [59], with the parameters allowed to vary with Gaussian constraints applied according to the corresponding predictions. The tauonic components are constrained to the branching fraction predictions in Ref. [59] with a relative width of 50%. Similar to Ref. [17], the modelling of this background is tuned and validated by fitting the one- and two-pion control regions. The former is used to determine the relative contribution of the lightest states (, , , and ) and the latter is used to determine the shape of heavier states such as the and mesons. As decays into the heavier states are particularly poorly known, a second-order polynomial-based shape variation is allowed to freely vary in the fit. The associated nuisance parameters are allowed to be different between the isolation regions in order to absorb the contribution of decays.
@@
The most signal-like background is that from decays since a muon originating from a charm hadron has similar kinematics to that originating from a decay. This background is modelled using simulation with shape corrections allowed to vary in the fit, for which the one-kaon control region is crucial. The fit model includes two distinct categories of this background. The first arises from two-body transitions, such as the decay , and the second arises from multi-body transitions, such as the decay . The two-body background is well reproduced by simulation whereas the multi-body modes are less well-understood, and a shape correction is therefore applied. The main correction consists of a pair of weights computed as linear and quadratic functions of the mass of the two charm hadrons. Corrections to the relative size of the and contributions to this background as well as the fraction of two-body to multi-body backgrounds are shared between all isolation regions. Unlike Ref. [17], the shape correction obtained in the one-kaon region is not applied directly to the signal region as a significant difference is observed in the parameters obtained from the two regions. This is not surprising as the Dalitz distribution of decays depends on the total charge of the combination, and the isolation requirements preferentially select certain charge combinations due to correlations with the presence of a meson. A separate correction based on the same parameterisation is applied for the signal region to allow for differences in the two regions.
@@
A complication is that the misidentification sample consists of several different particle types, each with a different probability to be misidentified as a muon. In order to estimate the relative mixture of particle types, the misidentification sample is split into regions based on the PID response. The criteria for each region is designed such that the region is dominated by a particular particle type. For example, an electron misidentification region is determined by requiring that the particle track is identified as an electron based on the calorimeter and RICH systems. Cross-feed between the regions is determined by fitting the distribution of the number of candidates in each region. Templates for each particle type are determined from dedicated control samples in data with the exception of fake tracks, which are obtained from simulation. Once the relative abundance of each particle type is determined as a function of the track kinematics, the control samples are used to extrapolate the background into the signal region. Furthermore, a smearing of the track momentum [17] is applied to account for the decay in flight of misidentified muons, which has a small but statistically significant impact on the template shape.
@@
![Refer to caption](x2.png)
![Refer to caption](x3.png)
![Refer to caption](x4.png)
![Refer to caption](x5.png)
@@ 6 Systematic uncertainties
Approximately 2% of the selected events contain multiple candidates, leading to a systematic which is determined by randomly selecting one candidate in those events and repeating the fit.
Finally, systematic uncertainties that were found to be negligible include the potential contribution from and decays, the determination of the neutral isolation selection efficiency, the assumptions behind the sPlot procedure and the effects of incomplete QED modelling in the simulation [60].
7 Results and conclusion
@@
where the first uncertainties are statistical and the second systematic. The correlation coefficient between the two measurements is . These results are from the SM predictions and 1.09 from the world average [9]. These are the first measurements of and using the decay mode at LHCb, the first analysis which uses tracker-only simulation and the first measurement to use HAMMER during the minimisation procedure of the likelihood fit. Assuming isospin symmetry between the charged and neutral decays, a combination with other LHCb measurements [16, 17] results in and , with a correlation coefficient of . Contributions to the systematic uncertainties due to sources common across the measurements are conservatively assumed to be 100% correlated. These contributions include the uncertainty associated with the form factors and branching fractions constrained from external measurements. The precision of these measurements is primarily limited by the size of the signal and control samples, so more precise measurements are expected with future LHCb datasets.

### v2

# EUROPEAN ORGANIZATION FOR NUCLEAR RESEARCH (CERN) âââ CERN-EP-2024-125

@@
LHCb collaborationâ â â Authors are listed at the end of this letter.
The branching fraction ratios of and decays are measured with respect to their muonic counterparts, using a data sample corresponding to an integrated luminosity of 2.0 fb-1 collected by the LHCb experiment in proton-proton collisions at . The reconstructed final states are formed by combining mesons with candidates, where the is reconstructed via the decay. The decay modes involving a meson are accessed via the decay , where the candidate is not reconstructed. The results are
@@
Published in Phys. Rev. Lett. 134 (2025) 061801
Š 2025 CERN for the benefit of the LHCb collaboration. CC BY 4.0 licence.
Â
@@
In the Standard Model (SM), the transition of a beauty quark () into a charm quark (), a tau lepton () and a tau anti-neutrino (), , proceeds via a tree-level boson exchange.111The inclusion of charge conjugate processes in this Letter is implied throughout. The SM decay amplitude is large, resulting in large branching fractions and copious signal yields. Despite this, transitions are highly sensitive to new physics as the process involves three third-generation fermions [1, 2, 3, 4]. This system therefore represents one of the best potential sources of contributions from new physics mediators that couple preferentially to the third generation [5, 6]. Furthermore, being a charged-current semileptonic transition, the decay amplitude can be factorised between the leptonic and hadronic parts to a good approximation. Uncertainties due to the strong interaction can be reduced by forming lepton universality ratios, where a transition is normalised to the corresponding muonic one, . In the SM, these ratios only deviate from unity due to the different lepton masses.
@@
The experimental challenge in analysing this system is the presence of multiple neutrinos in the final state, mandating a partial reconstruction of both the -lepton and -hadron decays. This partial reconstruction leaves the signal highly susceptible to a large number of backgrounds, which must be carefully controlled. Previous analyses of the system include measurements of the -lepton and charm-hadron polarisation [7, 8] and the lepton universality ratios and , which are formed from the branching fraction ratios . Lepton universality is the SM symmetry of equal coupling to the different charged lepton species, only broken by the different masses of the leptons. A combination [9] of the measurements [10, 11, 12, 13, 14, 15, 16, 17] differs by approximately three standard deviations from the SM predictions [18, 19, 20, 21, 22, 23, 24, 25, 26], indicating a potential enhancement of the semitauonic transition rate. Additional measurements of this process are therefore highly motivated.
@@
This Letter describes the first simultaneous measurement of the lepton universality ratios at the LHCb experiment, using the decays , and , specifically the first LHCb measurement of . This measurement uses data corresponding to 2.0 collected during the period 2015â2016 in 13âTeV collisions. Using isospin symmetry, these measurements can be combined with those of . The analysis strategy is similar to the existing measurement performed with the decay [17], where a three-dimensional template fit to the squared four-momentum transferred to the lepton pair (), the energy of the muon in the meson rest frame () and the squared invariant mass missing from the visible system () is performed to separate the signal and the normalisation from the backgrounds. The contribution from the first excited states ( and ) to the ground state is significantly reduced in this measurement compared to the case in Ref. [17]. The meson in the decay is not reconstructed, due to the large background from neutral objects.
@@
Simulated events are used to optimise the signal selection, model the signal and backgrounds, and calculate the relative efficiency between the signal and normalisation channels. In the simulation, collisions are generated using Pythia [30] with a specific LHCb configuration [32]. Decays of hadronic particles are described by EvtGen [33], in which final-state radiation is generated using Photos [34]. The interaction of the generated particles with the detector, and its response, are implemented using the Geant4 toolkit [35] as described in Ref. [37]. For this Letter, a new fast simulation technique is used whereby the photon propagation in the Ring Imaging CHerenkov (RICH) detectors, shower development in the calorimeters and activity in the muon stations are not simulated. This tracker-only simulation is a factor of eight faster and has a 40% smaller event size than full simulation. A smaller sample of simulated events which includes the full set of interactions (full simulation) is used for a cross-check and to develop specific tools such as those to reject partially reconstructed backgrounds with an additional neutral pion or photon in the final state.
@@
At the hardware stage, events are required to be triggered by activity unrelated to the signal candidate or by high transverse energy signals in the HCAL due to the decay products. In the software stage, one- or two-track combinations of the decay products are required to pass a multivariate selection based on the transverse momentum and impact parameter with respect to an associated primary vertex (PV). A subsequent software level selects a displaced three-track combination with a invariant mass compatible with the mass and large transverse momentum. The candidate is combined with a muon with a loose vertex-quality requirement. The muon is required to have a momentum above 3 to ensure that it reaches the muon stations. Particle identification (PID) requirements are applied on the final state particles, based on information from the RICH and muon systems.
@@
A set of offline selection critera, following the final trigger stage selection, is additionally applied. The candidate is required to have a significant impact parameter with respect to any PV, which reduces background from charm decays originating directly from the collision. A neural network classifier [38] is also applied to all final-state particles to reject track candidates which have more than 30% of hits that do not originate from the same source (fake tracks).
@@
Background for which the candidate is an accidental track combination from different sources is rejected by employing a Boosted Decision Tree (BDT) [39, 40] implemented in the TMVA toolkit [41]. The BDT is trained to distinguish simulated decays as a proxy for the semitauonic signal from background, which is modelled using data selected with invariant mass outside the range 1840â1900. The BDT features include kinematic and geometric information about the candidate as well as track quality variables for the decay products. The BDT requirement is optimised by maximising the metric , where is the expected number of signal decays from simulation and is the amount of background extrapolated from the mass sidebands in the data. This requirement is 97% efficient for the signal and rejects 35% of the background.
@@
Although the decay , where , is a signal channel in this analysis, the selection is optimised for the decay and therefore an isolation requirement is imposed based on neutral calorimeter objects, typically produced in partially-reconstructed background decays. A dedicated BDT is trained to discriminate between and decays, which is referred to as neutral isolation. This BDT uses the presence of neutral calorimeter objects in a cone around the direction and reconstructed candidates that have an invariant mass consistent with the known mass [43]. A requirement on the output of this BDT is used to remove 30% of candidates at the expense of 10% . The marginal gain is due to the large amount of activity in the ECAL from background processes.
@@
Finally, a BDT-based isolation tool for charged tracks [44], referred to as track isolation, is used to classify all additional tracks in the event based on whether they are consistent with originating from the signal -decay vertex. The three tracks that are most likely to originate from the signal candidate are used to control partially reconstructed backgrounds via a set of requirements on the BDT response that define isolation regions, described below.
@@
A maximum-likelihood fit to the mass distribution of the selected candidates is performed, where the signal is parameterised by an Hypatia function [45] and the background by an exponential function. The resulting yields of and background are 3.1M and 150K, respectively. The fit also determines per-event signal weights via the sPlot technique [46] that are used in the subsequent fit to the kinematic fit variables, , and , to subtract background where the combination does not originate from a meson.
@@
Several corrections are applied to improve the accuracy of the simulation. Firstly, the trigger and neutral isolation selection must be emulated for the tracker-only simulation. The hardware-level trigger must be emulated as it relies on information missing in the track-only simulation and is done by reproducing the two ways in which the events can be triggered. The efficiency for triggering via other particles in the event, which accounts for around 90% of the candidates, is measured using decays as a function of the kinematics. This is done directly using data in order to improve the description of the simulation. The emulation of the efficiency for one of the decay products to directly trigger through a high cluster in the HCAL is determined with fully simulated signal decays as a function of the kinematics of the decay products.
@@
The software trigger is accurately emulated by applying the same requirements offline, with the exception of the PID response, which is corrected for separately. The efficiency of the neutral isolation BDT requirement is modelled by applying a per-event weight, computed using a multivariate reweighter [47] trained on full simulation. With this approach, the efficiency can be parameterised as a function of the generator-level kinematic properties of the candidate, the sum of the four-momenta of the photons in the event, and the fit variables. The efficiencies of the PID requirements on the reconstructed final-state particles are determined using control samples [48] as a function of track kinematics and the number of tracks in each event. They are also used to determine misidentification probabilities, which are crucial to estimate the background where the muon candidate is a misidentified track. The tracking efficiency is determined using decays in data [49], as a function of the track kinematics. The kinematics of the meson as well as the event multiplicity are corrected using decays.
@@
Finally, a weight-based calibration procedure is designed to correct a comprehensive set of geometric and kinematic variables associated with the final-state particles as well as the Dalitz variables of the decay. In addition to the properties of the candidate, the kinematics of the three additional tracks found by the track-isolation algorithm are also corrected. This procedure is designed to correct for residual mismodelling of the simulation, such as the trigger response. The calibration sample for this correction procedure is selected candidates in the region , where the sample is enriched in the normalisation channels. The result of this procedure depends on the relative abundance of the different decays assumed in the simulation, which is estimated by an initial fit to the data sample. The corrected simulation is consistent with the data in the distributions of all considered variables.
@@
The corrected simulation is used to determine the relative efficiency between the signal and normalisation channels. This efficiency ratio is found to be 0.57 for the mode and 0.58 for the mode. These ratios are smaller than unity since the relatively low momentum muon from the lepton has a small detection efficiency, and the lifetime results in a poorly defined vertex. Systematic uncertainties on this ratio are subdominant compared to those affecting the signal yield determination.
@@
The tauonic signal decay is only 2% of the selected sample, split evenly between the and modes, meaning that control over the background is essential to obtain a precise and accurate measurement. A three-dimensional binned-likelihood fit to , and is therefore employed to fully exploit the kinematic differences in the signal and background channels. Each contribution to the fit is modelled using a template histogram derived either from simulation or from specifically selected control samples in data.
@@
Uncertainties in the modelling of the components are incorporated into the fit as bin-by-bin correlated variations in the templates. These variations are implemented via a polynomial interpolation [50]. Exceptions to this are variations in the form factors, which are performed using the HAMMER tool [51] as implemented in Ref. [52]. The HAMMER tool allows for a fast, but exact, variation of a template as a function of the decay-model parameters. This variation is important to allow the form factors of both the signal and normalisation channels to vary as the constraints derived from precise lattice calculations [53, 26, 54, 55] can have larger uncertainties than those obtained from the fit. The form factors for the and modes are parameterised according to the BGL formalism [56] with multivariate Gaussian constraints applied according to the predictions reported in Refs. [57, 26].
@@
The background originating from decays is less well-understood compared to those involving the lower mass states, despite a recent measurement by Belle [58]. This is because the measurements are less precise and the narrow-width assumption is less well-motivated due to the significant natural width of the states. The form-factor models employed for the decays are those described in Ref. [59], with the parameters allowed to vary with Gaussian constraints applied according to the corresponding predictions. The tauonic components are constrained to the branching fraction predictions in Ref. [59] with a relative width of 50% to the prediction values. Similar to Ref. [17], the modelling of this background is tuned and validated by fitting the one- and two-pion control regions. The former is used to determine the relative contribution of the lightest states (, , , and ) and the latter is used to determine the shape of heavier states such as the and mesons. As decays into the heavier states are particularly poorly known, a shape variation that depends linearly on the true of the decay is allowed to freely vary in the fit. The associated nuisance parameters are allowed to be different between the isolation regions in order to absorb the contribution of decays.
@@
The most signal-like background is that from decays, where and represent any number of particles originating from the or hadrons, since a muon originating from a charm hadron has similar kinematics to that originating from a decay. This background is modelled using simulation with shape corrections allowed to vary in the fit, for which the one-kaon control region is crucial. The fit model includes two distinct categories of this background. The first arises from two-body transitions, such as the decay , and the second arises from multi-body transitions, such as the decay . The two-body background is well reproduced by simulation whereas the multi-body modes are less well-understood, and a shape correction is therefore applied. The main correction consists of a pair of weights computed as linear and quadratic functions of the mass of the two charm hadrons, i.e. an empirical model similar to Ref. [17]. Corrections to the relative size of the and contributions to this background as well as the fraction of two-body to multi-body backgrounds are shared between all isolation regions. After the correction, a good fit quality in the one-kaon region is obtained, validating the procedure.
Unlike Ref. [17], the shape correction obtained in the one-kaon region is not applied directly to the signal region as a significant difference is observed in the parameters obtained from the two regions. This is not surprising as the Dalitz distribution of decays depends on the total charge of the combination, and the isolation requirements preferentially select certain charge combinations due to correlations with the presence of a meson. A separate correction based on the same parameterisation is applied for the signal region to allow for differences in the two regions.
@@
A complication is that the misidentification sample consists of several different particle types, each with a different probability to be misidentified as a muon. In order to estimate the relative mixture of particle types, the misidentification sample is split into regions based on the PID response. The criteria for each region is designed such that the region is dominated by a particular particle type. For example, an electron misidentification region is determined by requiring that the particle track is identified as an electron based on the calorimeter and RICH systems. Cross-feed between the regions is determined by fitting the distribution of the number of candidates in each region. Templates for each particle type are determined from dedicated control samples in data with the exception of fake tracks, which are obtained from simulation. Once the relative abundance of each particle type is determined as a function of the track kinematics, the control samples are used to extrapolate the background into the signal region. Furthermore, a smearing of the track momentum [17] is applied to account for the decay in flight of pions and kaons into muons, which has a small but statistically significant impact on the template shape.
@@
Approximately 2% of the selected events contain multiple candidates, leading to a systematic uncertainty, which is determined by randomly selecting one candidate in those events and repeating the fit.
Finally, systematic uncertainties that were found to be negligible include the potential contribution from and decays, the determination of the neutral isolation selection efficiency, the assumption that the reconstructed mass is uncorrelated with the fit variables in the sPlot procedure and the effects of incomplete QED modelling in the simulation [60].
@@
where the first uncertainties are statistical and the second systematic. The correlation coefficient between the two measurements is â. These results are from the average [9] of SM predictions [18, 19, 20, 21, 22, 23, 24], and , and 1.09 from the world average [9]. These are the first measurements of and using the decay mode at LHCb, the first analysis which uses tracker-only simulation and the first measurement to use HAMMER during the minimisation procedure of the likelihood fit. Assuming isospin symmetry between the charged and neutral decays, a combination with other LHCb measurements [16, 17] results in and , with a correlation coefficient of . This combination is also consistent with the SM prediction. Contributions to the systematic uncertainties due to sources common across the measurements are conservatively assumed to be 100% correlated. These contributions include the uncertainty associated with the form factors and branching fractions constrained from external measurements. The precision of these measurements is primarily limited by the size of the signal and control samples, so more precise measurements are expected with future LHCb datasets.